# Systematic Classification of History-Dependent Invariants in the Lorenz System: Regularization Classes and Dynamical Signatures

### v1

@@ VI.1 The Discrete $\mathbb{Z}_{2}$ Symmetry
VI.2 Parity Classification of Invariants
The invariants can be classified by the transformation properties of their polynomial parts $P_{n}(x,y,z)$ under $\mathcal{S}$. The bilinear terms transform as:
$x^{2},y^{2},z^{2},xy,xyz\to\text{(even)},$
$xz,yz\to\text{(odd)}.$ (69)
@@ III $K_{13}$, $K_{14}$ E
VI.3 Implications for Triad Classification
The canonical Triad $\mathcal{T}=\{K_{5},K_{11},K_{13}\}$ consists entirely of even invariants. Therefore
$P_{n}(\mathcal{S}(x,y,z))=P_{n}(x,y,z)$ (70)
for $n\in\{5,11,13\}$. Symmetric initial conditions produce identical fingerprints:
$(c_{1},c_{2},c_{3})\big|_{(x_{0},y_{0},z_{0})}=(c_{1},c_{2},c_{3})\big|_{(-x_{0},-y_{0},z_{0})}.$ (71)
The fingerprint therefore characterizes trajectories up to $\mathbb{Z}_{2}$ symmetry. Symmetric orbits share fingerprints, reflecting the fact that they are dynamically equivalent.
Parity-breaking Triads. To distinguish symmetric trajectories, one can employ an odd invariant. The Triad $\mathcal{T}^{\prime}=\{K_{3},K_{11},K_{13}\}$ includes the odd invariant $K_{3}$, yielding $(c_{1}^{\prime},c_{2},c_{3})|_{\mathcal{S}(\gamma)}=(-c_{1}^{\prime},c_{2},c_{3})|_{\gamma}$.
VI.4 Regularization Class and Sensitivity to Lobe Transitions
The Lorenz attractor possesses topological structure encoded in its template [6]. Periodic orbits can be classified by their symbolic dynamics: the sequence of left (L) and right (R) lobe visits.
The Class III polynomial parts $P_{\mathrm{III}}$ contain terms involving $xy$. Since the product $xy$ changes sign at lobe-switching events (when the trajectory crosses $x=0$), $\dot{P}_{\mathrm{III}}$ varies rapidly at these transitions.
At the separatrix, $\dot{x}=\sigma(y-x)$, so at $x=0$: $\dot{x}=\sigma y$. The product $xy$ evolves from negative (before crossing) through zero (at crossing) to positive (after crossing) during a left-to-right transition. The derivative $d(xy)/dt|_{x=0}=\sigma y^{2}>0$ confirms monotonic increase through crossing.
Class III invariants are therefore sensitive to separatrix position. The evolution function $Q_{\mathrm{III}}=-\dot{P}_{\mathrm{III}}$ exhibits large magnitude at lobe crossings. Class I polynomial parts also contain $xy$ terms but include $z^{2}$ contributions that distribute variance across the entire trajectory, reducing the relative prominence of separatrix spikes (Sec. VII.4).
@@ VII.1 High-Precision Verification
We performed numerical integration using 30-digit extended precision arithmetic in Mathematica with the StiffnessSwitching method, which automatically adapts to handle stiff intervals during separatrix crossings.
@@
    Conservation: $|K_{n}(T)-K_{n}(0)|<10^{-25}$ over $T=5000$ time units.
@@
    Zero mean: $|\langle Q_{n}\rangle_{T}|<10^{-10}$ for $T>1000$.
Table 3 presents conservation errors for the canonical Triad after one UPO period. Errors of order $10^{-36}$$10^{-38}$ confirm conservation to machine precision.
Table 3: Conservation errors |K(T)K(0)||K(T)-K(0)| for Triad invariants after one UPO period ($T\approx 1.56$).
@@ VII.2 Statistical Analysis of Auxiliary Evolution Functions
represents the instantaneous rate at which $v_{n}$ evolves to maintain conservation.
Pre-chaotic regime ($\rho=23$). Both classes yield statistically indistinguishable distributions (Fig. 2a). Kurtosis values $\kappa_{\mathrm{I}}\approx 36.8$ and $\kappa_{\mathrm{III}}\approx 37.0$ are effectively identical. This baseline establishes that there is no intrinsic bias in the Class III formulation.
Chaotic regime ($\rho=28$). A clear structural divergence emerges (Fig. 2b). Standardized distributions reveal different tail structures:
Differential intermittency. Class III exhibits higher kurtosis ($\kappa_{\mathrm{III}}\approx 15.1$) compared to Class I ($\kappa_{\mathrm{I}}\approx 7.9$), a 91% increase. The heavy tails indicate that Class III conservation is subject to sporadic large-amplitude events (“spikes”) punctuating quiescent intervals.
Asymmetry structure. Class III displays stronger negative skewness ($S_{\mathrm{III}}\approx-2.67$) versus Class I ($S_{\mathrm{I}}\approx-1.64$). Negative skewness indicates that extreme negative excursions are more intense than positive fluctuations. At separatrix crossings in the upper region of the attractor ($z>38$), $\dot{P}_{\mathrm{III}}>0$, yielding $Q_{\mathrm{III}}=-\dot{P}_{\mathrm{III}}<0$. The enhanced negative skewness reflects the concentration of large negative spikes at lobe switches.
Interpretation. The elevated kurtosis correlates with lobe-switching events. At separatrix crossings, $P_{\mathrm{III}}$ varies rapidly due to its $xy$-terms, forcing rapid adjustments in the Class III evolution function. Class I quantities lack this geometric sensitivity.
These dynamical signatures establish that the regularization class structure reflects genuine differences arising from polynomial structure, not algebraic convention.
Table 4: Statistical characterization of evolution functions $Q_{n}(t)$ for Class I ($K_{5}$) and Class III ($K_{13}$). Kurtosis $\kappa$ measures tail weight (Gaussian: $\kappa=3$); skewness $S$ measures asymmetry. Statistics computed from $N=9\times 10^{4}$ samples over $T=1000$ time units.
@@
Kurtosis $\kappa$ 36.8 37.0 7.9 15.1
Skewness $S$ 0.33-0.33 0.38-0.38 1.64-1.64 2.67-2.67
Pre-chaotic versus chaotic kurtosis. A noteworthy feature of Table 4 is that the pre-chaotic regime ($\rho=23$) exhibits substantially higher kurtosis ($\kappa\approx 37$) than the fully chaotic regime ($\kappa_{\mathrm{III}}\approx 15$). This counterintuitive result reflects a fundamental difference in the statistical structure of the two dynamical regimes. At $\rho=23$, the system operates near the subcritical Hopf bifurcation where trajectories exhibit complex transient dynamics: long intervals of quiescence near local attractors or unstable periodic orbits are punctuated by abrupt escapes. These rare but intense bursts generate distributions with pronounced heavy tails, yielding elevated kurtosis. In contrast, the fully developed chaos at $\rho=28$ produces more ergodic and mixing dynamics, with fluctuations distributed more uniformly across time. The increased frequency of moderate-amplitude events dilutes the statistical prominence of extreme excursions, thereby reducing kurtosis. Additionally, Class I and Class III distributions are statistically indistinguishable at $\rho=23$ ($\kappa\approx 37$ for both), while they diverge significantly at $\rho=28$ ($\kappa_{\mathrm{I}}=7.9$ vs. $\kappa_{\mathrm{III}}=15.1$). This divergence confirms that the class-dependent dynamical signatures emerge specifically from the interaction between polynomial structure and chaotic dynamics; they are not artifacts of the invariant construction itself.
VII.3 Algebraic Mechanism of Class III Spikes
The statistical analysis reveals that Class III invariants exhibit significantly higher intermittency in the chaotic regime. We identify the algebraic mechanism responsible.
The evolution function $Q_{n}(x,y,z)=-\dot{P}_{n}$ is polynomial in $(x,y,z)$; it contains no singular factors. The spikes in $Q_{\mathrm{III}}(t)$ arise not from singularities but from the rapid variation of $\dot{P}_{\mathrm{III}}$ as the trajectory traverses specific regions of phase space.
Two surfaces must be carefully distinguished:
  1. 1.
    Regularization surface (z-nullcline): $p_{\mathrm{III}}=xy-\beta z=0$, where the original auxiliary variable $u_{n}$ diverges.
  2. 2.
    Separatrix: the stable manifold of the saddle point at the origin, locally approximated by $x=0$.
These surfaces are generically distinct. The regularization procedure ensures that all invariants are smooth at the z-nullcline. The spikes arise from the rapid variation of $P_{\mathrm{III}}$ at the separatrix.
The Class III polynomial parts contain terms proportional to $xy$. The evolution function therefore includes:
$\frac{d(xy)}{dt}=\dot{x}y+x\dot{y}=\sigma(y-x)y+x[x(\rho-z)-y].$ (73)
At the separatrix $x=0$: $d(xy)/dt=\sigma y^{2}>0$. The product $xy$ evolves from negative through zero to positive during a left-to-right lobe transition.
Sign of spikes. For $K_{13}$ with $P_{13}=y^{2}/2+xyz-xy\rho+x^{2}\sigma/2-xy\sigma$, the time derivative at $x=0$ is:
$\dot{P}_{13}\big|_{x=0}=y^{2}(-1+\sigma z-\sigma\rho-\sigma^{2}).$ (74)
With $\sigma=10$ and $\rho=28$, this yields $\dot{P}_{13}|_{x=0}=y^{2}(10z-381)$. The sign depends on the altitude $z$ at which the separatrix crossing occurs: for $z>38.1$, $\dot{P}_{13}>0$ and hence $Q_{13}<0$; for $z<38.1$, the signs reverse. In the Lorenz attractor, lobe-switching events occur predominantly in the upper region where $z\approx 40$4545. Taking $z\approx 43$ as representative:
$\dot{P}_{13}\big|_{x=0}\approx y^{2}(430-381)=+49\,y^{2}>0.$ (75)
Therefore $Q_{13}=-\dot{P}_{13}\approx-49\,y^{2}<0$, confirming negative spikes (Fig. 2c). The negative skewness $S_{\mathrm{III}}\approx-2.67$ reflects this predominance of negative excursions at topological transitions.
Why Class I exhibits lower intermittency. The Class I invariant $K_{5}$ also contains an $xy$ term ($xy\rho$), yet exhibits lower kurtosis. The resolution lies in the polynomial structure: $P_{5}=y^{2}/2+\beta z^{2}/2-xy\rho$ contains the term $\beta z^{2}/2$, which $P_{13}$ lacks. The $z^{2}$ term creates continuous baseline fluctuations: since $z$ oscillates between approximately 5 and 45 on the attractor, the $z^{2}$ contribution varies by a factor of approximately 80, generating substantial variance in $Q_{5}$ even during quiescent intervals.
The statistical mechanism: kurtosis scales as $\kappa\propto\mu_{4}/(\sigma_{\mathrm{spike}}^{2}+\sigma_{\mathrm{base}}^{2})^{2}$. Adding baseline variance dilutes the prominence of spikes. Class I has larger total variance from the $z^{2}$-driven baseline, reducing kurtosis.
The Class III polynomial $P_{13}$ couples exclusively to $xy$-dependent terms. All significant variation concentrates at separatrix crossings. Between crossings, $Q_{13}$ remains quiescent. This concentration produces the heavy-tailed distributions characteristic of high kurtosis.
The elevated intermittency of Class III thus reflects not the mere presence of $xy$ terms (which Class I also possesses), but the absence of smoothing terms like $z^{2}$ that distribute variance across the trajectory.
The evolution function
$Q_{\mathrm{III}}(t)=-\dot{P}_{\mathrm{III}}(x(t),y(t),z(t))$ (76)
evaluated along trajectories inherits the rapid variations. The spike magnitude scales as:
$|Q_{\mathrm{III}}|_{\text{max}}\sim\frac{\Delta P_{\mathrm{III}}}{\tau_{\text{cross}}}.$ (77)
Numerical analysis confirms $|Q_{\mathrm{III}}|$ reaches $10^{4}$$10^{5}$ during lobe-switching events.
Distinction from singularities. The regularization procedure ensures that $K_{\mathrm{III}}$ is well-defined throughout phase space. The spikes occur at the separatrix ($x=0$), which is generically not the z-nullcline. The evolution function $Q_{\mathrm{III}}$ is smooth everywhere; the “spikes” are large values of a smooth function, not singularities. The high intermittency reflects geometric sensitivity to topological transitions, not numerical artifacts.
VII.4 Class III Spikes as Symbolic Markers
The Class III evolution function $Q_{\mathrm{III}}(t)$ functions as a symbolic marker that signals each addition to the trajectory’s symbolic sequence.
@@ VII.4 Class III Spikes as Symbolic Markers
In the standard symbolic dynamics description [6], each visit to the left lobe is encoded as $L$, each visit to the right lobe as $R$. The sequence $\ldots LLRLLRLR\ldots$ captures the trajectory’s itinerary through the attractor. Each lobe-switching event appends a new symbol to this sequence.
@@
Figure 2: Statistical signatures of the regularization class structure. Probability density functions of standardized evolution functions $Q_{\mathrm{I}}$ (Class I, solid blue) and $Q_{\mathrm{III}}$ (Class III, dashed red) on logarithmic scale. (a) Pre-chaotic regime ($\rho=23$): distributions are statistically indistinguishable ($\kappa\approx 37$ for both). (b) Chaotic regime ($\rho=28$): Class III develops heavier tails ($\kappa_{\mathrm{III}}\approx 15$ versus $\kappa_{\mathrm{I}}\approx 8$), indicating enhanced intermittency. (c) Temporal dynamics in the chaotic regime: $x(t)$ (blue) and $Q_{\mathrm{III}}(t)/5000$ (red, rescaled for visibility). Negative spikes in $Q_{\mathrm{III}}$ correlate with zero-crossings of $x(t)$ (vertical dashed lines marking separatrix crossings). The spikes arise from rapid variation of xyxy-terms at the separatrix; the invariant $K_{\mathrm{III}}$ remains constant throughout.
@@ VIII.1 Principal Result: Class-Dependent Dynamical Sensitivity
Class I invariants as continuous probes. The polynomial parts $P_{\mathrm{I}}$ contain terms like $z^{2}$ that vary continuously throughout the attractor. The evolution function $Q_{\mathrm{I}}=\dot{v}_{\mathrm{I}}$ exhibits moderate intermittency ($\kappa_{\mathrm{I}}\approx 8$), with variance distributed across quiescent and active intervals alike. These invariants track smooth, continuous aspects of the flow.
Class III invariants as topological probes. The polynomial parts $P_{\mathrm{III}}$ couple directly to the bilinear product $xy$, which changes sign at every lobe transition. The evolution function $Q_{\mathrm{III}}$ exhibits large negative spikes ($\kappa_{\mathrm{III}}\approx 15$) precisely at separatrix crossings, with quiescent intervals between transitions. These invariants function as geometric markers of the symbolic dynamics, encoding the discrete topological structure of the attractor.
This distinction is the primary result of the statistical analysis: the regularization class structure, which emerges from algebraic properties of the orthogonality ansatz, maps onto a meaningful partition of dynamical sensitivity. The elevated kurtosis of Class III is a direct consequence of polynomial structure concentrating significant variation into discrete spike events at lobe switches.

### v2

@@ VII.1 High-Precision Verification
We performed numerical integration using 80-digit extended precision arithmetic in Mathematica with the StiffnessSwitching method, which automatically adapts to handle stiff intervals during separatrix crossings. Initial conditions and parameters were specified using exact arithmetic (integers and rationals) to prevent precision contamination.
@@
    Conservation: $|K_{n}(T)-K_{n}(0)|<10^{-35}$ over $T=5000$ time units.
@@
    Zero mean: $|\langle Q_{n}\rangle_{T}|\to 0$ as $T^{-1}$.
Table 3 presents conservation errors for the canonical Triad after one UPO period. Errors of order $10^{-36}$$10^{-38}$ confirm conservation to numerical precision. With 80-digit working precision, the machine epsilon is $\sim 10^{-80}$; the observed errors of $O(10^{-37})$ reflect normal accumulation of roundoff during integration, confirming that the conservation laws are satisfied to the limits of numerical precision.
Table 3: Conservation errors |K(T)K(0)||K(T)-K(0)| for Triad invariants after one UPO period ($T\approx 1.56$). Calculations performed with 80-digit working precision.
@@ VII.2 Statistical Analysis of Auxiliary Evolution Functions
represents the instantaneous rate at which $v_{n}$ evolves to maintain conservation. Statistics were computed from $N=4\times 10^{6}$ trajectory samples (integration time $T=400$ after transient removal, sampling interval $\Delta t=10^{-4}$).
Pre-chaotic regime ($\rho=23$). Both classes yield statistically similar distributions (Fig. 2a). Kurtosis values $\kappa_{\mathrm{I}}\approx 28.1$ and $\kappa_{\mathrm{III}}\approx 28.3$ are effectively identical. This baseline establishes that there is no intrinsic bias in the Class III formulation; the statistical equivalence confirms that both classes respond identically to the simple spiral dynamics without lobe-switching.
Chaotic regime ($\rho=28$). A clear structural divergence emerges (Fig. 2b). The standardized distributions reveal different tail structures:
Differential intermittency. Class III exhibits substantially higher kurtosis ($\kappa_{\mathrm{III}}\approx 15.8$) compared to Class I ($\kappa_{\mathrm{I}}\approx 8.2$), a 93% increase. The heavy tails extending beyond $\pm 5\sigma$ indicate that Class III conservation is subject to sporadic large-amplitude events (“spikes”) punctuating quiescent intervals.
Asymmetry structure. Both classes display pronounced negative skewness in the chaotic regime, with Class III exhibiting even stronger asymmetry ($S_{\mathrm{III}}\approx-2.73$) than Class I ($S_{\mathrm{I}}\approx-1.66$). This indicates that extreme negative excursions dominate the fluctuation statistics for both classes, but the effect is amplified in Class III due to the rapid sign changes of the $xy$ term during separatrix crossings. The enhanced negative skewness of Class III reflects the geometric asymmetry of lobe-switching events: trajectories accelerate sharply as they approach the separatrix from one direction.
Interpretation. The elevated kurtosis of Class III correlates with lobe-switching events. At separatrix crossings, $P_{\mathrm{III}}$ varies rapidly due to its $xy$-terms, forcing rapid adjustments in the Class III evolution function. Class I quantities lack this geometric sensitivity due to the smoothing effect of the $z^{2}$ term in $P_{5}=y^{2}/2+\beta z^{2}/2-xy\rho$.
Pre-chaotic versus chaotic kurtosis. A noteworthy feature is that the pre-chaotic regime ($\rho=23$) exhibits substantially higher kurtosis ($\kappa\approx 28.2$) than the chaotic regime ($\kappa_{\mathrm{I}}\approx 8.2$). This counterintuitive result reflects a fundamental difference in statistical structure. At $\rho=23$, the system operates near the subcritical Hopf bifurcation where trajectories exhibit complex transient dynamics: long intervals of quiescence near local attractors are punctuated by abrupt escapes. These rare but intense bursts generate distributions with pronounced heavy tails, yielding elevated kurtosis. In contrast, fully developed chaos at $\rho=28$ produces more ergodic and mixing dynamics, with fluctuations distributed more uniformly across time.
Table 4: Statistical characterization of evolution functions $Q_{n}(t)$ for Class I ($K_{5}$) and Class III ($K_{13}$). Kurtosis $\kappa$ measures tail weight (Gaussian: $\kappa=3$); skewness $S$ measures asymmetry. Statistics computed from $N=4\times 10^{6}$ samples over $T=400$ time units.
@@
Kurtosis $\kappa$ 28.1 28.3 8.2 15.8
Skewness $S$ 0.16-0.16 0.44-0.44 1.66-1.66 2.73-2.73
VII.3 Class III as a Topological Probe
@@
The Class III evolution function $Q_{\mathrm{III}}(t)$ functions as a symbolic marker that signals each addition to the trajectory’s symbolic sequence. In the standard symbolic dynamics description [6], each visit to the left lobe is encoded as $L$, each visit to the right lobe as $R$. The sequence $\ldots LLRLLRLR\ldots$ captures the trajectory’s itinerary through the attractor.
@@
VII.4 Topological Precursor Detection
The differential response between Class III and Class I evolution functions provides the basis for a topological precursor detector. We construct the differential signal
$\Delta S(t)=Q_{13}(t)-Q_{5}(t),$ (53)
which isolates the topological sensitivity of Class III from the continuous dynamics tracked by Class I.
Detection protocol. Define an alert threshold $\mathcal{A}_{\mathrm{th}}$ based on the amplitude of $\Delta S(t)$. When $|\Delta S(t)|$ exceeds $\mathcal{A}_{\mathrm{th}}$, a precursor event is registered. The latency $\Delta t$ is defined as the time interval between the precursor signal and the subsequent separatrix crossing (zero-crossing of $x(t)$).
Scaling law. Figure 2(c) reveals a remarkable empirical relationship between the spike amplitude $\mathcal{A}$ and the prediction latency $\Delta t$. The data are well-described by a shifted power law:
$\Delta t=t_{\min}+C\mathcal{A}^{-n},$ (54)
with fitted parameters:
$t_{\min}$ $\approx 0.14,$ (55)
$n$ $\approx 1.0.$ (56)
Connection to the stable eigenvalue. Parametric analysis across the range $\beta\in[2.0,3.2]$ reveals that the exponent $n$ scales linearly with the stable eigenvalue: $n=m\beta+b$ with slope $m=0.196\pm 0.02$ and $b\approx 0.5$, with $R^{2}=0.94$. At the standard value $\beta=8/3$, this yields $n\approx 1.0$. The linear relationship is physically significant: the Class III regularization polynomial $p_{\mathrm{III}}=xy-\beta z$ explicitly contains $\beta$, and the predictive capability is governed by relaxation along the stable manifold (the $z$-direction near the fixed points). The proportionality constant $m\approx 0.2$ indicates that the differential sensor effectively integrates trajectory information over approximately five relaxation timescales ($\tau=1/\beta$) before generating a detectable precursor signal.
The physical interpretation is as follows. As the trajectory approaches the separatrix, the product $xy$ decreases toward zero while $z$ relaxes toward the unstable manifold. The rate of approach is governed by the stable eigenvalue $\beta$. Trajectories with larger initial departure from the separatrix (smaller $\mathcal{A}$) require longer relaxation times before crossing, yielding the inverse power-law relationship.
Intrinsic latency. The parameter $t_{\min}\approx 0.14$ represents the minimum prediction horizon, determined by the phase-space separation between the unstable manifold and the Poincaré section $x=0$. Even for the strongest precursor signals, the trajectory requires a finite time to traverse this geometric gap.
Bimodal structure. The scatter plot in Fig. 2(c) reveals a secondary “phantom branch” at low amplitudes and long latencies (inset). These events correspond to trajectories entering the alert zone far from the actual separatrix crossing, subsequently executing partial orbits before the true transition. The bimodal distribution directly manifests the heavy-tailed statistics documented in Table 4.
Figure 2: Statistical signatures and topological precursor detection. Probability density functions of standardized evolution functions $Q_{\mathrm{I}}$ (Class I, solid blue) and $Q_{\mathrm{III}}$ (Class III, dashed red) on semi-logarithmic scale, with Z-score normalization. (a) Pre-chaotic regime ($\rho=23$): distributions are statistically similar ($\kappa\approx 28.2$ for both), establishing that no intrinsic bias exists in the Class III formulation. (b) Chaotic regime ($\rho=28$): Class III develops heavier tails ($\kappa_{\mathrm{III}}\approx 15.8$ versus $\kappa_{\mathrm{I}}\approx 8.2$) extending beyond $\pm 5\sigma$, indicating enhanced intermittency at topological transitions. Both classes exhibit pronounced negative skewness ($S_{\mathrm{I}}\approx-1.66$, $S_{\mathrm{III}}\approx-2.73$). (c) Scaling law for the topological precursor. Latency $\Delta t$ (time from precursor signal to separatrix crossing) versus spike amplitude $\mathcal{A}$ of the differential signal $\Delta S=Q_{13}-Q_{5}$. Gray points: individual events; black circles: logarithmically binned averages; green curve: fitted shifted power law $\Delta t=t_{\min}+C\mathcal{A}^{-n}$ with $t_{\min}\approx 0.14$ and $n\approx 1.0$ at $\beta=8/3$. Parametric analysis reveals that $n$ scales linearly with $\beta$ (slope $m\approx 0.196$, $R^{2}=0.94$), establishing that predictive capability is governed by relaxation along the stable manifold. Statistics computed from $N=4\times 10^{6}$ trajectory samples.
@@ VIII.1 Principal Result: Class-Dependent Dynamical Sensitivity
Class I invariants as continuous probes. The polynomial parts $P_{\mathrm{I}}$ contain terms like $z^{2}$ that vary continuously throughout the attractor. The evolution function $Q_{\mathrm{I}}=\dot{v}_{\mathrm{I}}$ exhibits moderate intermittency ($\kappa_{\mathrm{I}}\approx 8.2$), with variance distributed across quiescent and active intervals alike. These invariants track smooth, continuous aspects of the flow.
Class III invariants as topological probes. The polynomial parts $P_{\mathrm{III}}$ couple directly to the bilinear product $xy$, which changes sign at every lobe transition. The evolution function $Q_{\mathrm{III}}$ exhibits elevated kurtosis ($\kappa_{\mathrm{III}}\approx 15.8$) with spikes synchronized to separatrix crossings and quiescent intervals between transitions. These invariants function as geometric markers of the symbolic dynamics, encoding the discrete topological structure of the attractor.
Differential detection. The complementarity between classes enables construction of a differential detector $\Delta S=Q_{\mathrm{III}}-Q_{\mathrm{I}}$ that acts as a “topological band-pass filter”: it cancels the continuous rotational dynamics to isolate the hyperbolic singularities associated with lobe-switching events.